OpenStudy (keana):

Choose two trinomials from the list below to factor. Using complete sentences, explain how to factor each one. 2x2 + 13x + 15 2x2 + 27x + 13 5x2 - 2x - 7 10x2 - 7x - 3 8x2 - 11x + 3

OpenStudy (anonymous):

2x2 + 13x + 15 =2x2+10x+3x+15=2x(x+5) +3(x+5)=(2x+3)(x+5)

OpenStudy (anonymous):

2x2 + 27x + 13 =2x2 + 26x+x+13=2x(x+13)+1(x+13)=(2x+1)(x+13)